OpenStudy (lacypennelll):

what is the solution to g/5 = 2/3

OpenStudy (abb0t):

mutiply 5 to both sides.

OpenStudy (aravindg):

MULTIPLY BY 5 ON BOTH SIDES

OpenStudy (lacypennelll):

o.o....

OpenStudy (aravindg):

lol why the scary face?

OpenStudy (aravindg):

@lacypennelll got it ?

OpenStudy (lacypennelll):

yea ^~^

OpenStudy (aravindg):

great ;)

OpenStudy (anonymous):

g/5 = 2/3 Well, here.. g=2x5/3 (Cross MULTIPLYING) as we need the value of "g", we have to cross multiply.. g=10/3 so.. the answer is 3.33

OpenStudy (anonymous):

mutiply 5 to both sides.?? how? can anybody tell me that?

OpenStudy (lacypennelll):

Lai here let me explain

OpenStudy (anonymous):

yeah!! sure!

OpenStudy (lacypennelll):

brb

OpenStudy (lacypennelll):

Cross multiply g/5 x 3/2

OpenStudy (lacypennelll):

so it would be g x 2 = 3 x 5

OpenStudy (lacypennelll):

2g 15 -- = -- 2 2

OpenStudy (lacypennelll):

:P mmk

OpenStudy (lacypennelll):

The answer wou be 7.5 in the end

OpenStudy (anonymous):

Oh, man! i don't need the answer.. it's your question.. well. okay.. you have put so much effort for explaining me.. so lemme give u a medal! thanks!

OpenStudy (lacypennelll):

Lol thanks

OpenStudy (aravindg):

7.5 ??

OpenStudy (aravindg):

10/3=3.33

OpenStudy (lacypennelll):

Well I got 7.5 and I got it right when I submitted it..

OpenStudy (anonymous):

yeah!! i was right. huh!! ^_^

OpenStudy (aravindg):

What??

OpenStudy (lacypennelll):

O_O wow my test is jack then

OpenStudy (aravindg):

Where did you submit it?

OpenStudy (aravindg):

I am 200% sure its 3.33

OpenStudy (lacypennelll):

FLVS In my end of the module test O_O

OpenStudy (aravindg):

ok do you understand laiba's working @lacypennelll ?

OpenStudy (lacypennelll):

Yea

OpenStudy (anonymous):

thanks @AravindG

OpenStudy (aravindg):

yw :)